OpenStudy (anonymous):

Which of the following functions is a solution to the differential equation y" + 2y'

OpenStudy (zzr0ck3r):

take their first and second derivatives and sub them in for y'', and y'

OpenStudy (anonymous):

What do you mean? What am I supposed to take the derivative of?

OpenStudy (zzr0ck3r):

they should have gave you functions

OpenStudy (zzr0ck3r):

i.e. there is not enough information to solve this as you asked it..

OpenStudy (anonymous):

a) y=e^t b)-e^t c) y=e^-t d) e^-2t

OpenStudy (anonymous):

Okay so first do I take the derivatives of each of these?

OpenStudy (zzr0ck3r):

so for a) \(y'=e^t, y''=e^t\) now plug them into the thing you have and see if makes since \(y''+2y'+y=0\implies e^t+2e^t+e^t=0 \) for all \(t\), this is not true, so its not a solution.

OpenStudy (anonymous):

The derivative of -e^t is 0

OpenStudy (zzr0ck3r):

no \((-e^t)'=-e^t\)

OpenStudy (anonymous):

-e^t+2(-e^t)+-e^t

OpenStudy (zzr0ck3r):

and so is the second derivative, so \(y''+2y'+y=-e^t+2(-e^t)+(-e^t)=-4e^t\ne 0\)

OpenStudy (zzr0ck3r):

try the next one:)

OpenStudy (anonymous):

anything that is e raised to the power has the same derivative right?

OpenStudy (zzr0ck3r):

no, only \(e^x\) is its own derivative

OpenStudy (anonymous):

e^-t +2(e^-t) +(e^-t)=

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

=4e^-t

OpenStudy (zzr0ck3r):

\(y=e^{-t}\\y'=-e^{-t}\\y''=-(-e^{-t})=e^{-t}\) plugging this in \(y''+2y'+y=e^{-t}+2(-e^{-t})+e^{-t}=2e^{-t}-2e^{-t}=0\)

OpenStudy (anonymous):

One minute I am trying to write all of this down so that I can try and do it by myself again. Haha. okay so for some reason I though that e^anything was always the same

OpenStudy (anonymous):

But I guess it is only e^x. Right?

OpenStudy (zzr0ck3r):

no, there is the chain rule \(y=e^{ax}\implies\frac{dy}{dx}=e^{ax}(ax)'=e^{ax}a=ae^{ax}\)

OpenStudy (zzr0ck3r):

so correct the only two things that will ever look the same after you take the derivative is \(\large y= e^x \text{ and } y=-e^x\)

OpenStudy (zzr0ck3r):

but not \(\large e^{-x}\)

OpenStudy (anonymous):

oh okay that makes sense